OpenStudy (anonymous):

find the real or imaginary solutions of each equation by factoring x^4-8x^2=-16

OpenStudy (paounn):

write the expression as \[x^4 -8x^2 +16 =0\] Recognize in it the square of \[(x^2 -4)\] Your equation becomes \[[(x+2)(x-2)]^2 = 0 \], whoose soultions should be evident.

OpenStudy (anonymous):

is the solutions (2i,2i,-2i,-2i)

OpenStudy (anonymous):

is that right

OpenStudy (paounn):

\[(2i)^4 -8 (2i)^2 = -16 \rightarrow 16*1 - 8 * 4 *(-1) = -16 \rightarrow 16 + 32 = -16 \rightarrow 48 = -16\] which is evidently false. Same goes for -2i

OpenStudy (anonymous):

so wats the answer im completely confused

OpenStudy (paounn):

From my first post, remembering that \[(a*b)^2 = a*b*a*b\]\[(x+2)(x-2)(x+2)(x-2)=0\] so you get that \[x=-2, x=2, x=-2, x=2\]

OpenStudy (anonymous):

i figured it our actually before u wrote back but thank u